# Deepfake Generation and Detection: A Benchmark and Survey

### v1

@@ 3.1.3 Talking Face Generation
In this section, we review current methods from three perspectives: audio/text driven, multimodal driven, and diffusion-based. We also summarize them in Tab. IV.
$\bullet$ Audio/Text Driven. Methods aim to map and guide lip and facial movements in generated videos by understanding the semantic information from the driving source [255, 256, 237, 242]. Early methods [117, 257] perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers [103, 118, 104] propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk [258] decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. AD-NeRF [236] directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic neural radiance field, synthesizing highly realistic facial videos. SadTalker [240] extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods [259, 95, 243, 260] propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos [261, 239]. Recently, some methods [262, 263, 238, 245] extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN [262] establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO [263] employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP [264] introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et al[241] propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 [19] explores practical strategies for reducing the training workload.
@@
TABLE V: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.).
@@ 3.1.4 Facial Attribute Editing
$\bullet$ Irrelevant Attributes Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images [271, 275, 21]. GuidedStyle [275] leverages attention mechanisms in StyleGAN [5] for the adaptive selection of style modifications for different image layers. IA-FaceS [21] embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches [274, 29, 28, 278] combine GANs with NeRF [26] for enhanced spatial awareness capabilities. Specifically, FENeRF [274] uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ [278] enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
@@
TABLE VI: Overview of representative foreign detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD. The full table can be viewed in Appendix B.
@@
Space Domain Gram-Net [286] CVPR’20 ➆➈ ➆➈ The method posits that genuine faces and fake faces exhibit inconsistencies in texture details.
Zhao et al[288] CVPR’21 ➀➁➂ A texture enhancement module, an attention generation module, and a bilinear attention pooling mod -ule are proposed to focus on texture details.
@@
Ciamarra et al[294] WACV’24 By analyzing the depicted surface features in an image, the inconsistency of the unambiguous environ -ment represented by pixel values is exploited as a detection criterion.
Time Domain Yang et al[94] ICASSP’19 Focusing on the inconsistency in the head pose in videos by comparing the estimated head pose using all facial landmarks with the one estimated using only the landmarks in the central region.
FTCN [295] ICCV’21 ➀➁➂➃[180] It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
@@
Data-Driven Yu et al[298] ICCV’19 [69] [69] Assuming that GAN models generate images with a unique fingerprint, targeted forgery detection during training can be achieved.
@@
3.2 Foreign Detection
In this section, we review current foreign detection techniques based on the type of detection cues, categorizing them into three: Space Domain (Sec. 3.2.1), Time Domain (Sec. 3.2.2), and Data-Driven (Sec. 3.2.3). We also summarize the detailed information about popular methods in Tab. VI.
@@ 3.2.1 Space Domain
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color [31], saturation [303], artifacts [288, 290, 291], gradient variations [30], etc. Specifically, Zhao et al[288] propose a detection method based on a multi-head attention mechanism, framing deep forgery detection as a fine-grained image classification problem. RECCE [291] considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad [30] utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions [289], as well as the fine-grained details of image textures [304, 286].
@@
$\bullet$ Frequency Inconsistency. Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works [36, 306, 37, 307] utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, F3-Net [36] proposes a dual-branch framework. One frequency-aware branch utilizes Frequency-aware Image Decomposition (FAD) to learn subtle forgery patterns in suspicious images. In contrast, the other branch aims to extract high-level semantics from Local Frequency Statistics (LFS) to describe the frequency-aware statistical differences between real and forged faces. HFI-Net [307] consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration [308, 292], and Guo et al[292] design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
@@ 3.2.2 Time Domain
$\bullet$ Inter-Frame Inconsistency. Methods [295, 297, 33, 34] based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et al[297] focus on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et al[33] design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et al[34] approach DeepFake detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames.
@@
3.2.3 Data-Driven
@@ 4.1 Main Results on Deepfake Generation
TABLE VIII: Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ MPIE [328], ➍ ETH-Xgaze, ➎ Gaze360 [224], ➏ MPIIGaze [225], ➐ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
@@
Free-HeadGAN [223] ➊➍➎➏ 0.810 22.16 0.100 35.40 -
HiDe-NeRF [227] ➊➋➑ 0.931 21.90 0.084 - 0.862
@@
TABLE X: Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRw, ➍ CelebV-HQ, ➎ TalkingHead-1KH, ➏ MPIE  [328].
@@ 5 Future Prospects
$\bullet$ Foreign Detection. With the rapid development of facial forgery techniques, the central challenge in face forgery detection technology is accurately identifying various forgery methods using a single detection model. Simultaneously, ensuring that the model exhibits robustness when detecting forgeries in the presence of disturbances such as compression is crucial. Most detection models follow a generic approach targeting common operational steps of a specific forgery method, such as the integration phase in face swapping or assessing temporal inconsistencies, but this manner limits the model’s generalization capabilities. Moreover, as forgery techniques evolve, forged videos may evade detection by introducing interference during the detection process.
@@ B. Supplementary Tables
This section supplements some of the table content in the main text, further expanding the number of models introduced. Tab. XVI supplements Tab. II in the main text, and Tab. XVII supplements Tab. VI in the main text.
@@
Method Year Dataset Categorize Limitation Highlight
Traditional Graphics
Blanz et al[171] EG’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
@@
ReliableSwap [195] arXiv’23 ➊➋➌➍ Other Poor ability to preserve face feature attributes. Constructing a supervisor called the "cyclic triplet" enhances the model’s identity preservation capability.
@@
TABLE XVII: Overview of representative foreign detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD.
@@
Frank et al[306] ICML’20 [69] Severe artifacts in GAN-generated images can be easily identified in the frequency space, which stem from the upsampling steps used in GANs.
@@
HFI-Net [307] TIFS’22 ➁➂➃➅[77] It is noted that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
@@
Guo et al[292] TIFS’23 ➀➁ ➀➁➂ Constructing a new backbone network for Deepfake detection by designing air-frequency interaction convolution.

### v3

@@ 3.1.3 Talking Face Generation
In this section, we review current methods from three perspectives: audio/text driven, multimodal conditioned, diffusion-based, and 3D-model Technologies. We also summarize them in Tab. IV.
$\bullet$ Audio/Text Driven. Methods aim to map and guide lip and facial movements in generated videos by understanding the semantic information from the driving source [259, 255, 243]. Early methods [119, 105] perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers [120, 106] propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk [238] decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. SadTalker [241] extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods [260, 97, 261, 262] propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos [263, 240]. Recently, some methods [264, 265, 266] extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN [264] establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO [265] employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP [267] introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et al[268] propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 [19] explores practical strategies for reducing the training workload.
@@
$\bullet$ 3D-model Technologies. 3D models, exemplified by NeRF, are gaining traction in talking face generation [254, 255, 256]. AD-NeRF [254] directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic NeRF. AE-NeRF [256] employs a dual NeRF framework to separately model audio-related regions and audio-independent regions. Furthermore, some methods [257, 258] adopt Tri-Plane [271] to represent facial and audio attributes. Synctalk [257] models and renders head motion using a tri-plane hash representation, and then further enhances the output quality using a portrait synchronization generator. In the foreseeable future, 3D Gaussian Splatting [272] will also be widely applied to this task.
TABLE V: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.). Full table can be viewed in Appendix B.
@@
SDGAN et al[287] AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
FaceDNeRF [288] NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.
@@ 3.1.4 Facial Attribute Editing
$\bullet$ Irrelevant-attribute Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images [276, 291, 21]. GuidedStyle [291] leverages attention mechanisms in StyleGAN [5] for the adaptive selection of style modifications for different image layers. IA-FaceS [21] embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches [279, 29, 28, 282] combine GANs with NeRF [26] for enhanced spatial awareness capabilities. Specifically, FENeRF [279] uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ [282] enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
@@
TABLE VI: Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD. Full table can be viewed in Appendix B.
@@
Space Domain Face X-ray [294] CVPR’20 ➀➁➂➉ Focusing on boundary artifacts of face fusion for forgery detection.
@@
Ba et al[299] AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on information bottleneck theory, is proposed for constraint.
Time Domain FTCN [300] ICCV’21 ➀➁➂➃[182] It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
@@
Choi et al[303] CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
Xu et al[304] IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
Frequency Domain F3-Net [36] ECCV’20 A two-branch frequency perception framework with a cross-attention module is proposed.
HFI-Net [305] TIFS’22 ➁➂➃➅[77] Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
Tan et al[307] AAAI’24 ➀➄ A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.
Data Driven Dang et al[308] CVPR’20 ➂➅ Utilizing attention mechanisms to handle the feature maps of the detection model.
@@
Finfer [310] AAAI’22 ➀➂➇[92] Based on an autoregressive model, using the facial representation of the current frame to predict the facial representation of future frames.
@@
3.2 Forgery Detection
In this section, we review current forgery detection techniques based on the type of detection cues, categorizing them into three: Space Domain (Sec. 3.2.1), Time Domain (Sec. 3.2.2), Frequency Domain (Sec. 3.2.3), and Data-Driven (Sec. 3.2.4). We also summarize the detailed information about popular methods in Tab. VI.
@@ 3.2.1 Space Domain
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color [31], saturation [312], artifacts [313, 296, 297], gradient variations [30], etc. Specifically, RECCE [297] considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad [30] utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions [295], as well as the fine-grained details of image textures [314, 315]. Recently, Ba et al[299] focus not only on the discordance in a single image region but also on the detection of fused local representation information from multiple non-overlapping areas.
@@ 3.2.2 Time Domain
$\bullet$ Inter-Frame Inconsistency. Methods [300, 302, 33, 34, 303, 304] based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et al[302] focus on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et al[33] design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et al[34] approach DeepFake detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames. Choi et al[303] discover that the style variables in each frame of Deepfake work change. Based on this, they developed a style attention module to focus on the inconsistency of the style latent variables between frames.
@@
3.2.3 Frequency Domain
Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works [36, 320, 37, 305, 307] utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, HFI-Net [305] consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Tan et al[307] introduce a novel frequency-aware approach called FreqNet, which focuses on the high-frequency information of images and combines it with a frequency-domain learning module to learn source-independent features. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration [321, 306], and Guo et al[306] design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
3.2.4 Data Driven
@@ 4.1 Main Results on Deepfake Generation
TABLE VIII: Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 [226], ➎ MPIIGaze [227], ➏ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
@@
Free-HeadGAN [225] ➊➌➍➎ 0.810 22.16 0.100 35.40 -
HiDe-NeRF [230] ➊➋➐ 0.931 21.90 0.084 - 0.862
@@
TABLE X: Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRW, ➍ CelebV-HQ, ➎ TalkingHead-1KH.
@@ 5 Future Prospects
$\bullet$ Forgery Detection. With the rapid development of facial forgery techniques, the central challenge in face forgery detection technology is accurately identifying various forgery methods using a single detection model. Simultaneously, ensuring that the model exhibits robustness when detecting forgeries in the presence of disturbances such as compression is crucial. Most detection models follow a generic approach targeting common operational steps of a specific forgery method, such as the integration phase in face swapping or assessing temporal inconsistencies, but this manner limits the model’s generalization capabilities. Moreover, as forgery techniques evolve, forged videos may evade detection by introducing interference during the detection process.
@@ B. Supplementary Tables
This section supplements some of the table content in the main text, further expanding the number of models introduced. Tab. XVI supplements Tab. 2 in the main text, Tab. XVII supplements Tab. 4 in the main text, Tab. XVIII supplements Tab. 5 and Tab. XIX supplements Tab. 6 in the main text.
@@
Method Venue Dataset Categorize Limitation Highlight
Traditional Graphics Blanz et al[173] EG’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
@@
TABLE XVII: An overview table of talking face generation methods. Corresponding labels for the commonly used datasets are ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ.
Method Year Dataset Limitation Highlight
Audio / Text - Driven Chen et al[105] ECCV’18 [348][349] Poor resolution, inability to control pose and emotional. Proposed a novel generator network and a comprehensive model with four complementary losses, as well as a new audio-visual related loss function to guide video generation.
Zhou et al[350] AAAI’19 [343] Inability to control pose and emotional variations. Generate high-quality talking face videos by disentangling audio-visual representations.
Chen et al[120] CVPR’19 [348] Inability to control pose and emotional variations. Proposed a cascaded approach, using facial landmarks as an intermediate high-level representation.
MakeItTalk [238] TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
Ji et al[239] CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
SPACE [266] ICCV’23 ➋➌[351] Lack emotional and other latent attributes control Constructed a novel facial intermediate representation, achieving control over head pose, blinking, and gaze direction.
SadTalker [241] CVPR’23 ➏➒ Lack emotional and other latent attributes control. Based on the idea of 3DMM and conditional VAE, 3D coefficients controlling facial motion and expre -ssion are generated from audio to realize the reproduction of accurate face from audio.
Wang et al[249] TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
DREAM-Talk [352] arXiv’23 ➌➏ High model complexity. Proposed a two-stage approach to effectively capture subtle emotional nuances and ensure accurate lip synchronization.
3D-Model AD-NeRF [254] ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
EmoTalk [242] ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
Ye et al[258] ICLR’24 [76] Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.
TABLE XVIII: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.).
Method Venue Categorize Dataset Highlight
AttGAN [353] TIP’19 G. Applying attribute classification constraints to generated images has validated the drawbacks of enforcing stringent attribute independence constraints in latent representations.
TransEditor [290] CVPR’22 G. ➌➊ Emphasizing dual-space GAN interaction’s importance, a transformer architecture is introduced for improved interaction.
SDGAN et al[287] AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
FaceDNeRF [288] NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.
TABLE XIX: Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD.
@@
Ba et al[299] AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on inf -ormation bottleneck theory, is proposed for constraint.
Tian et al[358] TIFS’24 ➀➁➂➇ ➀➁➂➇ A weakly supervised model using a classification attention region module for forgery clue detection and a complementary learning module for en -hanced clue learning is introduced.
@@
Choi et al[303] CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
Xu et al[304] IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
Peng et al[361] TIFS’24 ➀➂➇ [92] ➀➂➇ [92] Adopting the idea of physiological signal detection, focusing on the temporal and spatial gaze inconsistency in video face forgery detection.
Frequency Frank et al[320] ICML’20 [69] Severe artifacts in GAN-generated images can be easily identified in the frequency space, which stem from the upsampling steps used in GANs.
HFI-Net [305] TIFS’22 ➁➂➃➅[77] Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
Tan et al[307] AAAI’24 ➀➄[362] A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.